OpenStudy (babe2323433):

What are the zeros of the polynomial function f(x) = x3 − x2 − 12x? a. 0,-4,-3 b. 0,-4,3 c. 0,4,-3 d. 0,4,3

OpenStudy (babe2323433):

Will fan and metal!!!

OpenStudy (jdoe0001):

hint: \(\bf f(x)=x^3-x^2-12x\iff x(x^2-x-12) \\ \quad \\ hint:\qquad -4\cdot 3=12\qquad \qquad -4+3=-1\) use those hints, and we'll give you a medal :)

OpenStudy (babe2323433):

Can you walk me through the problem?

OpenStudy (jdoe0001):

hmmm -12 I meant of course \(\bf f(x)=x^3-x^2-12x\iff x(x^2-x-12) \\ \quad \\ hint:\qquad -4\cdot 3=-12\qquad \qquad -4+3=-1\)

OpenStudy (jdoe0001):

-4 and 3 are factors of -12, the constant in the trinomial use those factors to construct two binomials then use FOIL to see if those two binomials give you the original trinomial :)

OpenStudy (babe2323433):

Im confused..

OpenStudy (babe2323433):

@sunnnystrong

OpenStudy (sunnnystrong):

@babe2323433 .. Okay so basically you need to factor this out. If \[ f(x) = x^3 − x^2 − 12x\] You can factor out an x & solve \[ f(x) = x(x^2 − x − 12)\] Now... you need to factor the quadractic: What two #'s when added together =-1 & when multiplied = -12 ---> gonna be -4 & 3 So: \[ f(x) = x(x-4)(x+3)\] Zeros of f are when x=3 so: set each factor = 0 and solve for x you get: x=0 , x=4, x=-3 are the roots of f

OpenStudy (babe2323433):

Ohhhhhhhhhhhhh.. Omg thank you sooo much!!!!

OpenStudy (sunnnystrong):

NP Always happy to help lol :D

OpenStudy (babe2323433):

How do I give you a metal? @sunnnystrong

OpenStudy (jdoe0001):

\(\begin{array}{llll} x({\color{blue}{ x^2-x-12 }} ) \end{array}\implies \begin{array}{cccllll} x^2&-x&-12\\ &-4+3&-4\cdot 3 \end{array} \\ \quad \\ thus\implies x^2-x-12\iff (x-4)(x+3) \\ \quad \\ x({\color{blue}{ x^2-x-12 }} )\iff x{\color{blue}{ (x-4)(x+3) }} \\ \quad \\ \textit{to get the zeros, set them to 0} \\ \quad \\ 0=x(x-4)(x+3)\implies \begin{cases} 0=x&\to 0=x\\ 0=x-4&\to 4=x\\ 0=x+3&\to -3=x \end{cases}\)

OpenStudy (sunnnystrong):

You click on the little metal lol

OpenStudy (babe2323433):

Ohh.. Lol Can you help me with a few more?

OpenStudy (sunnnystrong):

yep just close and the question and post another one :D

OpenStudy (jdoe0001):

yeah, you click on the "metal" to give it just make sure is not rusty, otherwise you may give the other fellow a rash now for the "medal", just click on "best response" button on the right :)